OpenStudy (anonymous):

What is the value of g[f(4)] for the functions f(x) = 2x + 1 and g(x) = 2x - 5? (Type only the value of your final answer, no variables.)

OpenStudy (lalaly):

f(4)=2(4)+1 = 8+1 =9 g(f(4))=g(9) = 2(9)-5=18-5=13

OpenStudy (anonymous):

first one is 9. second one is 3.

OpenStudy (anonymous):

lalaly y did u say 13?

OpenStudy (anonymous):

its 13 my bad.

OpenStudy (lalaly):

because its 13 2(9)=18 18-5=13

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

thanks everyone! i need more math help so if u can please help please do!